OpenStudy (anonymous):

Adult tickets for a play cost $9 and child tickets cost $4. If there were 27 people at a performance and the theatre collected $143 from ticket sales, how many children attended the play?

OpenStudy (anonymous):

11 children and 11 adults. 9x + 4x = 143 13x = 143 x=11

OpenStudy (anonymous):

Adult tickets for a play cost $9 and child tickets cost $4. If there were 27 people at a performance and the theatre collected $143 from ticket sales, how many children attended the play? A. 19 children B. 21 children C. 20 children D. 7 children

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

its D

OpenStudy (whpalmer4):

a = # of adults c = # of children a+c = 27 9a + 4c = 143 Solve by substitution or elimination

OpenStudy (anonymous):

Thank you walidfarhat it's confusing to me the way it is worded.

OpenStudy (luigi0210):

It'd be easier to substitute

OpenStudy (anonymous):

listen to luigi

OpenStudy (anonymous):

what i did was try everything its D 7x4=28+9(15)= 143

OpenStudy (luigi0210):

so either set the top equal to a or c, and plug it into the other equation

OpenStudy (whpalmer4):

Elimination is easy, too: multiply first eqn by -4: \[-4a -4c = -108\]\[9a+4c = 143\] --------------- \(5a=35\) etc.

OpenStudy (whpalmer4):

It's not D. D is 7 children and 27-7 = 20 adults. Adults pay $9 each, so right there that's $180, which is > the $143 collected. Always, always check your result to make sure it actually solves the problem when doing word problems (or any problems, really).

OpenStudy (anonymous):

Thank you all for your help